OpenStudy (anonymous):

3x^3-x^2+9x-3=0 this equation has 3 roots, how can we find them by using the quadratic formula?

hero (hero):

Why do you need the quadratic formula to find them?

OpenStudy (anonymous):

well...do you have a better way to find the three roots? i want to try it with the quadratic formula

hero (hero):

You don't use the quadratic formula for a cubic

OpenStudy (anonymous):

we can, it will be like 3x^2-x+9=0

hero (hero):

3x^3-x^2+9x-3=0 x^2(3x - 1) + 3(3x - 1) = 0 (3x - 1)(x^2 + 3) = 0 3x - 1 = 0 x^2 + 3 = 0 You can use quadratic formula to figure out the roots for x^2 + 3 = 0

OpenStudy (anonymous):

so the roots are 3 rooted, and 1/3... wheres the third? to the original equation

hero (hero):

x = -sqrt(3) x = +sqrt(3) x = 1/3